# KVzap: Fast, Adaptive, and Faithful KV Cache Pruning

### v1

@@ 1 Introduction
Pioneered by $H_{2}O$ (h2o), the Awesome-KV-Cache-Compression repository now lists dozens of KV cache pruning methods, with 20+ implemented in NVIDIA/kvpress. While these methods validate the pruning intuition—e.g., KVzip (kvzip) can reach $4\times$ compression with no accuracy loss on some tasks—none have been integrated into major inference engines such as vLLM (vllm), SGLang (sglang), or TRT-LLM (trtllm). In retrospect, each solution fails to meet at least one of the following criteria:
@@
  • We introduce KVzap, a new KV cache pruning technique which applies these surrogate models to the hidden states to prune KV pairs below a fixed threshold $\tau$.
@@ 2.3 KVzap
To address KVzip’s limitations, we train a per-layer surrogate—either a linear layer or a two-layer MLP—to predict $H$ scores $\log(s^{+})$ directly from the input hidden states $\mathbf{h}$ (we use log-space to match the exponential nature of softmax attention). The model acts independently at each sequence position $t$: it maps $\mathbf{h}_{t}\in\mathbb{R}^{D_{h}}$ to scores in $\mathbb{R}^{H}$, where $D_{h}$ is the hidden dimension and $H$ is the number of KV heads. Because it uses only one or two matrix multiplications and depends only on hidden states, KVzap is computationally efficient (see Appendix B) and can be applied during decoding.
@@
Another key difference lies in the eviction policy. Whereas KVzip enforces a fixed budget (e.g., keeping exactly 50% of KV pairs), KVzap uses thresholding, discarding KV pairs whose predicted score falls below a fixed threshold $\tau$. Higher thresholds yield higher compression ratios. This makes KVzap input-adaptive: it dynamically adapts the compression rate based on the prompt information density, retaining more tokens for complex inputs and fewer for redundant ones.
@@
Algorithm 1: PyTorch pseudocode for KV cache pruning using KVzap during prefilling. Decoding is similar but needs a score buffer to enforce the sliding window
@@ 3.3 Prefilling and decoding tasks
KV cache pruning is most impactful for tasks involving thousands of tokens, during prefilling (long inputs) or decoding (long outputs). To assess KVzap across these regimes, we evaluate KVzap-Linear and KVzap-MLP on two long-context benchmarks—RULER (ruler) ($n=6500$) and LongBench (longbench) (n=4750n=4750)—and one reasoning benchmark, AIME25 (aime25) ($n=30$).
@@ 3.8 Ablations
KVzap uses score thresholding rather than fixed top-$k$ selection, allowing the compression rate to adapt to prompt complexity. Figure 5 (left) highlights this adaptability, showing up to 20% variation across prompts. As shown in Figure 5 (right), thresholding outperforms fixed-ratio top-$k$ selection, whether per-head or per-layer (AdaKV, (adakv)).
@@ 4 Discussion
First, while results on a 32B model are encouraging, further validation is needed on larger open-source models (e.g., GLM 4.7 (glm45), Qwen3-235B-A22B (qwen3)) and architectures with sparse attention (e.g., DeepSeek V3.2, (DeepSeekV32)). Evaluation could also be extended to more reasoning benchmarks, agentic tasks, and short-context knowledge tasks.

### v2

@@ 1 Introduction
Pioneered by $H_{2}O$ (h2o), the Awesome-KV-Cache-Compression repository now lists dozens of KV cache pruning methods, with 20+ implemented in NVIDIA/kvpress. Yet, none have been integrated into major inference engines such as vLLM (vllm), SGLang (sglang), or TRT-LLM (trtllm). We argue that a practical pruning method must meet the following criteria:
@@
  • We introduce KVzap, a new KV cache pruning technique which applies these surrogate models to the hidden states, adopting DMS-style (dms) inference: per-head compression, score-based pruning via thresholding, and delayed eviction with a sliding window.
2 Related work
Block-Sparse Attention and Offloading. To mitigate the quadratic complexity of attention, recent research has explored architectural modifications and block-wise computation. Sparse attention mechanisms, such as DSA in DeepSeek V3.2 (DeepSeekV32), employ a lightweight indexer that computes coarse-grained similarity scores to retrieve only the most relevant KV blocks, reducing computational complexity to near-linear but requiring architectural support during pre-training. MoBA (moba) partitions input sequences into independent segments, allowing only specific blocks to attend globally, which reduces prefilling latency but often requires fine-tuning to recover performance. SeerAttention (seerattention) integrates hierarchical compression and selection directly into training. While these methods improve throughput, they focus on skipping computation rather than permanently reducing the memory footprint of the resident cache.
Training-Free Eviction Methods. KV cache pruning methods try to identify KV pairs that are unlikely to be used by the attention mechanism and remove them, hence reducing memory while preserving the information needed for faithful generation. These methods typically apply eviction after attention, leaving the forward pass unchanged. Early approaches rely on heuristics to identify and evict less critical tokens without model retraining. $H_{2}O$ (h2o) uses accumulated attention scores to retain “heavy hitters”, and StreamingLLM (streamingllm) keep recent tokens via sliding windows. While pioneering, these heuristics degrade significantly even at modest compression ratios. A variety of methods have since emerged (snapkv; pyramidkv; scissorhands; fastgen), though most fail to meet all of our criteria (see Awesome-KV-Cache-Compression for a comprehensive list). Many of these methods allow variable per-head compression rates, including AdaKV (adakv) and Duo Attention (duoattention). Among these, KVzip (kvzip) achieves the best performance in our benchmark (Figure 1) by defining token importance via “context reconstruction”, retaining only the KV pairs necessary to reproduce the input. However, KVzip requires a computationally expensive double forward pass on an extended prompt, making it unsuitable for online inference (Criterion 1) and it cannot be applied during decoding (Criterion 2).
End-to-End Learned Eviction. The most comparable approach to ours is Dynamic Memory Sparsification (DMS) (dms), which also learns a policy to dynamically prune the KV cache. DMS learns per-token eviction decisions via trainable weights applied to hidden states, using Gumbel-sigmoid reparametrization for differentiable training. The model is trained end-to-end to preserve its output distribution while masking KV pairs, with a loss combining KL divergence and a sparsity penalty. In contrast to KVzap, DMS evicts tokens before the attention pass, resulting in sparse prefilling. KVzap further differs in its training objective: it trains surrogate MLPs to approximate KVzip scores rather than using end-to-end distillation. We were unable to include DMS in our benchmark comparison due to the lack of publicly available checkpoints.
Quantization. Orthogonal to token pruning, quantization methods such as KIVI (kivi) and ZipCache (zipcache) reduce KV cache memory by lowering numeric precision, demonstrating that KV pairs can be stored in INT4 or even INT2 formats with specialized kernels; these approaches can be combined with pruning for additional compression.
@@ 3.3 KVzap
To address KVzip’s limitations, we train a per-layer surrogate—either a linear layer or a two-layer MLP—to predict $H$ scores $\log(s^{+})$ directly from the input hidden states $\mathbf{h}$, as proposed in DMS (dms) (we use log-space to match the exponential nature of softmax attention). The model acts independently at each sequence position $t$: it maps $\mathbf{h}_{t}\in\mathbb{R}^{D_{h}}$ to scores in $\mathbb{R}^{H}$, where $D_{h}$ is the hidden dimension and $H$ is the number of KV heads. Because it uses only one or two matrix multiplications and depends only on hidden states, KVzap is computationally efficient (see Appendix B) and can be applied during decoding.
@@
Another key difference lies in the eviction policy. Whereas KVzip enforces a fixed budget (e.g., keeping exactly 50% of KV pairs), KVzap - like DMS - uses thresholding, discarding KV pairs whose predicted score falls below a fixed threshold $\tau$. Higher thresholds yield higher compression ratios. This makes KVzap input-adaptive: it dynamically adapts the compression rate based on the prompt information density, retaining more tokens for complex inputs and fewer for redundant ones.
@@
Algorithm 1: PyTorch pseudocode for KV cache pruning using KVzap during prefilling; similar to DMS (dms) but applied after attention. Decoding is similar but needs a score buffer to enforce the sliding window
@@ 4.3 Prefilling and decoding tasks
KV cache pruning is most impactful for tasks involving thousands of tokens, during prefilling (long inputs) or decoding (long outputs). To assess KVzap across these regimes, we evaluated KVzap-Linear and KVzap-MLP on two long-context benchmarks—RULER (ruler) ($n=6500$) and LongBench (longbench) (n=4750n=4750)—and one reasoning benchmark, AIME25 (aime25) ($n=30$).
@@ 4.8 Ablations
KVzap uses score thresholding rather than fixed top-$k$ selection, allowing the compression rate to adapt to prompt complexity. Figure 5 (left) highlights this adaptability, showing up to 20% variation across prompts. As shown in Figure 5 (right), thresholding outperforms fixed-ratio top-$k$ selection, whether per-head or per-layer (AdaKV (adakv)).
@@ 5 Discussion
First, while results on a 32B model are encouraging, further validation is needed on larger open-source models (e.g., GLM 4.7 (glm45), Qwen3-235B-A22B (qwen3)) and architectures with sparse attention (e.g., DeepSeek V3.2 (DeepSeekV32)). Evaluation could also be extended to more reasoning benchmarks, agentic tasks, and short-context knowledge tasks.